# Language Model Council: Democratically Benchmarking Foundation Models on Highly Subjective Tasks

### v2

@@ 4 Results and Findings
Table 1 shows the main results of our Council-based EI study, highlighting the following insights:
@@
Like other benchmarks, larger models within the same family tend to outrank their smaller or older versions. However, unlike other benchmarks, Qwen-1.5-110B (#20 on Chatbot Arena) scores highest on our EI task, followed by GPT-4o (#1 on Chatbot Arena) [^10] This was the ranking as of May 2024. The Qwen-1.5 models have been dropped from Chatbot Area since the Qwen-2.5 family of models have been added.. This is surprising, as Qwen-1.5-110B does not typically outperform GPT-4o. One possible reason for this outcome is the use of Qwen-1.5-32B as the reference model (Appendix D). Because all of Qwen-1.5-110B’s responses are compared to the responses of a strictly smaller variant of the same family, this could result in an outsized advantage for Qwen-1.5-110B in the evaluation overall. This raises an interesting implication for Arena-style evaluations, as it’s possible that the choice of reference model may introduce a bias favoring its successors in the same arena.
@@
Using chain-of-thought (CoT) prompting, judges provided detailed reasoning for their preferences. We analyzed 1,000 reasoning traces and identified common themes (Appendix G).
@@
Despite a suggested 250-word limit, some models generated much shorter responses, even though decoding parameters allowed for longer output. Models that adhered to the limit, using 220+ words on average, performed better, while models in the bottom four positions averaged less than 200 words. Notably, Gemini-1.5-pro placed last with an average response length of just 115 words, far worse than its predecessor Gemini-1.0-pro (4th place) with 228 words on average. LLM Judges bias towards longer responses, but excluding the outliers (models that went well under the limit) made length bias insignificant (Table 6).
@@
Consistent judging, neutral voting patterns, and lower contrarianism lead to higher separability.
Our analysis shows that judges with more consistent votes and neutral voting patterns tend to achieve higher separability. Judges with high conviction — those that express strong preferences (A>>B or B>>A) frequently — are negatively correlated with separability, suggesting that overly strong votes tend to introduce noise. For example, claude-3-opus, despite expressing strong preferences only 0.1% of the time, achieves the second-highest separability (72.6%). Higher contrarianism, which measures how often a judge disagrees with the majority decision, also correlates with lower separability. Qwen1.5-72B, the most contrarian judge, has only 38.9% separability.
LLMs are lightly self-biased, but it is mitigated in the full Council.
@@
Agreement among LMC members, humans, and between the LMC and humans is similar.
@@
Figure 4: Spearman correlations for different evaluation benchmark scores for 9 LLMs.
@@
The LMC achieves the highest correlation with human-established rankings, outperforming other benchmarks, including those from similar domains like the EI-specific subset of Chatbot Arena (Appendix F) and EQ-Bench Paech (2023) (Figure 4). The LMC outperforms most individual LLM judges in terms of correlation with human rankings. While a few individual LLM judges, like dbrx-instruct, achieve similar correlation scores (e.g., 0.917), they do so with significantly lower separability (50.5% vs. 90.5% for the LMC). This suggests that the LMC’s collective judgment not only aligns more closely with human preferences but also provides clearer distinctions between model performance, making it a more reliable method for ranking models in our case study.
@@
The 20x20 LLM interactions generate a wealth of data, some of which is not included in the main paper for brevity. Additional insights can be found in Appendix B.
@@ 5 Discussion
One of the most important questions about using a LMC is whether it is worth the trouble. Collecting more opinions costs more energy, and parsing, tallying, and storing them add logistical costs. What is the value of another opinion, much less everyone’s opinion?
From a cost perspective, a closely related question is how many examples we should collect judgements for to begin with. Using a smaller test set simplifies stakes for human reviewers — it is far easier to read and remember 10 examples than 10,000 – while also reducing costs.
If we assume that the main goal is to establish good relative rankings, we narrow the focus of quantifying the success of a council by measuring the significance and stability of the final ranking. For significance, we look at separability (as in the main experiment), which measures how well models can be distinguished based on non-overlapping confidence intervals. For stability, we create a new metric, Mean Expected Rank Variance (MERV), defined as the expected ordinal swing of the average respondent’s rank (Appendix H). A MERV of 3 means an average respondent’s rank is expected to change up to 3 positions in a new trial. Lower MERV indicates that the benchmark has more stable rankings, with MERV of 0 signifying perfect deterministic-like stability.
@@ 5.1 Monte Carlo simulations
To study the dynamics of separability and stability with differently-sized councils and test sets irrespective of the inclusion of any specific LLM, we use a Monte Carlo procedure to simulate many random hypothetical councils and test sets.
Our Monte Carlo simulation procedure is as follows: 1) For a given council size $c$ and test set size $t$, randomly sample $c$ LLMs to form a council $C$ and $t$ examples to form a test set $T$. Sampling is performed with replacement. 2) Find the associated judgements from the main experiment for the specific ($C$, $T$) configuration to determine the scores and relative rankings for all LLMs. 3) Repeat for 100 trials [^11] We use 100 to be consistent with the number of rounds of bootstrapping in the main experiment. For calculating separability in Monte Carlo simulations, the trials themselves are used to bootstrap confidence intervals directly.. 4) After all 100 trials are complete, tally the results: for stability, observe fluctuations in rankings for each LLM to compute MERV, and for significance, report the mean separability. Figure 5 shows results for a sweep $c\in\{1,3,5,\dots,19\},\quad t\in\{10,20,30,\dots,100\}$.
@@
Both separability and stability improve as the number of test examples and the number of LLM judges increase, with the best scores achieved when both are maximized. However, based on the gradient maps for MERV (Figure 5(b)) and separability (Figure 5(d)), the added benefit of including either an additional judge or more test examples diminishes significantly in a concentric shape that starts  50 examples and  9 judges. The gradients in this darker zone indicates where the utility of any additional opinion—whether in the form of new test data or another LLM judge—becomes marginal.
@@
As the size of the LMC grows to include many different LLMs, it may become difficult to verify the quality of every member, particularly on subjective tasks. In our Monte Carlo procedure, we experiment with a simulation setting with adversarial judges. An adversarial judge is a fake LLM council member that returns ratings at random [^12] There are other adversarial algorithms that can be used such as always voting for the first position that we did not explore.. In Figure 6, we find that on both separability and stability, larger councils reduce the negative impact of adversarial LLM judges. We find that this robustness continues to increase as the size of the council grows, even when maintaining the same ratio of adversarial judges to real judges, albeit with diminishing marginal utility.
@@
With respect to stability and separability, it depends. For small test sets, the value of additional test examples outweighs the benefit of more judges, but when test data is abundant test data, the value of an additional judge is higher. If there are more than 20-30 examples in the test set, adding an additional judge provides greater value, but for less, it is better to get opinions on an additional batch of examples with the current set of judge(s).
Overall, larger councils are more robust to adversarial judges so the bigger your council is, the less concerned you need to be about a particularly bad judge from wreaking havoc on your study, and you don’t have to be as picky about your council members.
@@ 5.2 Oligarchical councils
If a task does benefit from multiple perspectives and the evaluation budget allows for a limited number of opinions, whose opinions should we include? Verga et al. (2024) suggests replacing a single large judge with 3 smaller judges. [^13] Council composition is likely a task-specific optimization; On HotpotQA, an ensemble of LLM judges was worse than using Claude Haiku alone Verga et al. (2024). We explore a similar but reverse question: can a subset of judges (or a single judge) effectively represent the fully democratic LMC?
We compare the rankings of three hand-curated sub-councils (Table 4): flagships, smalls, and top-4. We refer to the sub-council as an “oligarchical” council since we rely on a small subset of participating LLM judges to determine the ranking of every LLM, akin to oligarchies in human society.
Our analysis finds that full council participation is neither necessary to achieve rankings that align with human rankings nor maintain good separability. Smalls, a council composed of the smallest LLMs, achieves a respectable separability of 71%, which is already better than the average judge (53.3%). It also achieves a Spearman correlation of 0.88 with human rankings, only slightly lower than the full council’s correlation of 0.92 (Figure 23).
Still, it’s worthwhile to be thoughtful about council composition. For instance, top-4 achieves the same correlation with human evaluations as smallest (0.88), but with notably higher separability (79%). While smaller councils consisting of smaller LLMs can perform reasonably well, there are other factors to consider like separability, robustness, human alignment, and bias mitigation.
@@ 6 Conclusion
In this paper, we introduce the Language Model Council (LMC), a flexible, decentralized evaluation framework for ranking LLM agents through democratic participation. Applying the LMC to an emotional intelligence task with 20 LLMs, we demonstrate that the LMC can produce highly separable rankings that align more closely with human judgments than other benchmarks or individual judges. Through both Monte Carlo simulations and hand-curated sub-councils, we find that while larger councils provide benefits, they are incrementally diminishing, and the majority of key qualities—such as ranking significance, stability, and alignment with human evaluations—can be achieved with a smaller, well-chosen ensemble of judges. As humans increasingly rely on LLMs to evaluate other LLMs, we hope the LMC framework, along with insights from our case study, offers a valuable foundation for developing reliable LLM evaluations, even for highly subjective tasks.
@@
LLMs are inherently stochastic, meaning the same model can produce different ratings even with temperature set to 0 Chann (2023). Reproducibility is further complicated by closed-weight models like GPT-4, which may receive undisclosed updates. All responses in our study were collected in May 2024, but serverless providers, like Together, may update their APIs or discontinue support for certain models, as happened with Qwen-1.5-32B (replaced by Qwen-2.5). For open source models, changes in deployment hardware or GPU configuration can introduce additional variability, making exact replication of results difficult.
@@
While the LMC mitigates self-enhancement and position bias, it remains vulnerable to systemic biases within the evaluation framework itself. For instance, we speculate that using a single reference model may introduce an advantage for successors of the reference model in the arena. This specific aspect could be mitigated by using multiple reference models, but it can be challenging for new test data to compensate for a bias in fundamental evaluation procedures.
@@ Acknowledgements
Justin led the research. Amanda and Flor led the human study. Benji ran Monte Carlo simulaitons and generated gradient charts. We thank Sahand Sabour for creating EmoBench and insightful discussions about emotionally rich synthetic data. We thank Alex Tamkin for proposing the idea to measure the value of the incremental judge. We thank Mitchell Gordon for his suggestion to qualitatively analyze why certain responses are preferred. We thank David So for his ideas on calibration, repeatability, and model affinity. We thank Federico Bianchi and Sam Paech for their idea of measuring oligarchical councils and for their feedback on prompt design. Sam Paech provided insightful discussions on separability, voting aggregation, length bias, and the relationship to other leaderboards, and he advised us throughout the project. Finally, we thank Predibase for their support and funding this research.
Appendix A Language Model Council Membership
United States Open AI gpt-4-0613 OpenAI et al. (2023) 06/23 1246 86.4 [Uncaptioned image] Proprietary
United States Meta llama-3-70b-chat-hf Dubey et al. (2024) 04/24 1208 82.0 70 B Llama 3 Community
United States Meta llama-3-8b-chat-hf Dubey et al. (2024) 04/24 1153 68.4 8 B Llama 3 Community
Canada Cohere command-r Cohere (2024a) 04/24 1147 68.2 35 B CC-BY-NC-4.0
China Alibaba qwen1.5-110B-chat Bai et al. (2023) 02/24 1164 80.2 100 B Qianwen LICENSE
China Alibaba qwen1.5-72B-chat Bai et al. (2023) 02/24 1152 77.4 72 B Qianwen LICENSE
China Alibaba qwen1.5-32B-chat Bai et al. (2023) 02/24 1126 74.3 32 B Qianwen LICENSE
LLM All Flagships Smalls Top-4
gpt-4-turbo-04-09 [Uncaptioned image] [Uncaptioned image]
gpt-4-0613 [Uncaptioned image]
gpt-3.5-turbo-0125 [Uncaptioned image]
mistral-large-latest [Uncaptioned image] [Uncaptioned image]
open-mixtral-8x22b [Uncaptioned image]
open-mixtral-8x7b [Uncaptioned image] [Uncaptioned image]
llama-3-70b-chat-hf [Uncaptioned image] [Uncaptioned image]
llama-3-8b-chat-hf [Uncaptioned image] [Uncaptioned image]
gemini-1.0-pro [Uncaptioned image] [Uncaptioned image]
dbrx [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
command-r-plus [Uncaptioned image] [Uncaptioned image]
command-r [Uncaptioned image] [Uncaptioned image]
claude-3-sonnet-20240229 [Uncaptioned image]
claude-3-haiku-20240307 [Uncaptioned image] [Uncaptioned image]
qwen1.5-110B-chat [Uncaptioned image] [Uncaptioned image]
qwen1.5-72B-chat [Uncaptioned image]
qwen1.5-32B-chat [Uncaptioned image] [Uncaptioned image]
Appendix B Additional Findings
@@
Can we proactively remove inconsistent votes?
Table 1 shows a wide variety of judging qualities within the council, especially in consistency and separability. In the two-game setup, we gather ratings for models in both positions during pairwise comparisons, allowing us to identify and potentially remove inconsistent ratings before calculating ELO scores.
Automatically removing inconsistent votes affects the weighting of LLM judges, as those with more inconsistent ratings will have fewer votes counted and thus less influence overall. However, this could be a positive outcome, as positional consistency is often a sign that the judgment was arbitrary, especially in cases where tie ratings are not permitted. By comparison, some arena-based evaluation systems like Zheng et al. (2024) allow judges to declare a tie between responses, however these battles are completely excluded from ELO scoring.
In our emotional intelligence (EI) case study, we chose to retain all inconsistent votes, allowing downstream processes like aggregation, Bradley-Terry scoring, and bootstrapping to implement any diminished influence caused by inconsistency.
For the hand-selected sub-councils analyzed in Section 5.2, Table 5 shows the changes in key judging qualities when considering only consistent votes. Table 6 shows the impact of inconsistent vote pre-filtering on length bias.
@@
Due to table size constraints, we have divided the full tables of key judging qualities into several smaller tables. The full-sized 20x20 heatmaps from our LMC EI case study also span multiple pages. For ease of reference, all large tables and figures are compiled here:
@@
Figure 10: Heatmap of the affinities of LLM judges to LLM respondents. The relatively consistent horizontal bands in the heatmap suggest a clear consensus on the preferred LLM participants. Some judges, like mistral-large, exhibit high polarization, with a significant difference between their highest and lowest-rated LLMs. In contrast, judges like llama3-8b display a narrower range of affinity. The highest affinity expressed by any LLM comes from mistral-large for qwen1.5-110B, while the strong horizontal blue band for gemini-1.5-pro indicates it was a consensus low performer.
@@
Figure 12: Graph of top 5 affinities. An edge exists from LLM $a$ to LLM $b$ if $\text{affinity}(a,b)$ is in the top 5 affinities for LLM $a$. This view allows us to identify popular LLMs, hipster LLMs, and "LLM friendships" (where two LLMs have strong mutual affinity for each other). Popular LLMs, such as GPT-4o, have many arrows pointing toward them, while unpopular LLMs like dbrx-instruct have none. Qwen-1.5-32B and command-r form a "friendship" with mutual strong affinities, though Qwen-1.5-32B also receives incoming edges from three other LLMs, making it more widely liked. Some LLMs have only one "fan," others have several, and some have none. Of course, this analysis is somewhat contrived, as affinities are continuous values and using the top 5 as a cutoff is arbitrary. Nevertheless, it offers an interesting starting point for studying the dynamics and patterns that emerge when an arbitrary affinity threshold is established.
@@
Appendix C LLM Judge Calibration
To understand the reliability and natural variability of LLM model judges, we collect pairwise preference ratings on three responses to the same interpersonal conflict using different temperatures and pairwise comparison options. Two responses are competitive, and one is intentionally generic to serve as a ranking baseline (Figure 17). We measure invariability and Pairwise Positional Consistency (PPC), defined below.
C.1 Invariability

### v4

@@ 4 Results and Findings
Table 1 presents the main results of our LMC EI case study, with key insights summarized below.
@@
Like other benchmarks, larger models within the same family tend to outrank their smaller or older versions. However, unlike other benchmarks, Qwen-1.5-110B (#20 on Chatbot Arena) scores highest on our EI task, followed by GPT-4o (#1 on Chatbot Arena). [^9] The scores and rankings referened for Chatbot Arena were those as of May 2024. The Qwen-1.5 models have been dropped from Chatbot Area since the Qwen-2.5 family of models have been added. This is surprising, as Qwen-1.5-110B does not typically outperform GPT-4o. One possible reason for this outcome is the use of Qwen-1.5-32B as the reference model (Appendix C). Because all of Qwen-1.5-110B’s responses are compared to the responses of a strictly smaller variant of the same family, this could result in an outsized advantage for Qwen-1.5-110B in the evaluation overall. This raises an interesting possibility of successor bias in arena-style evaluations: the choice of reference model may inadvertently favor its successors in the same arena.
@@
Using chain-of-thought (CoT) prompting, judges provided detailed reasoning for their preferences. We analyzed 1,000 reasoning traces and identified common themes (Appendix F).
@@
Despite a suggested 250-word limit, some models generated much shorter responses, even though decoding parameters allowed for longer output. Models that adhered to the limit, using 220+ words on average, performed better, while all of the models in the bottom four positions averaged less than 200 words. Notably, Gemini-1.5-pro placed last with an average response length of just 115 words, far worse than its predecessor Gemini-1.0-pro (4th place) with 228 words on average. LLM judges bias towards longer responses, but if we exclude the models that went well under the limit, length bias becomes insignificant (Table 4).
@@
Consistent judging, neutral voting patterns, and lower contrarianism correlate with higher separability.
Judges with more consistent votes and neutral voting patterns tend to achieve higher separability. Judges with high conviction — those that express strong preferences (A>>B or B>>A) frequently — are negatively correlated with separability, suggesting that overly strong votes tend to introduce noise. For example, claude-3-opus, despite expressing strong preferences only 0.1% of the time, achieves the second-highest separability (72.6%). Higher contrarianism, which measures how often a judge disagrees with the majority decision, also correlates with lower separability. Qwen-1.5-72B, the most contrarian judge, has only 38.9% separability.
LLMs exhibit slight self-bias, but this effect is mitigated within the full Council.
@@
Agreement among LMC members, among humans, and between the LMC and humans is similar.
@@
Figure 4: Kendall-Tau correlation between benchmark scores and human study scores for nine LLMs (see Appendix D).
@@
The LMC achieves the highest correlation with human-established rankings, outperforming other benchmarks, including those from similar domains like the EI-specific subset of Chatbot Arena (Appendix E) and EQ-Bench Paech (2023) (Figure 4). While a few individual LLM judges, like dbrx-instruct, achieve similar correlation scores (e.g., 0.917), they do so with significantly lower separability (50.5% vs. 90.5% for the LMC). This suggests that the LMC’s collective judgment not only aligns more closely with human preferences but also provides clearer distinctions between model performance, making it a more reliable method for ranking models in our case study.
@@
The 20x20 LLM interactions generate a wealth of data, some of which is not included in the main paper for brevity. Additional insights can be found in Appendix A.
@@ 5 Discussion
One of the most important questions about using a Language Model Council is whether it is worth the trouble. Collecting more opinions costs more energy, and parsing, tallying, and storing them add logistical costs. What is the value of another opinion, much less everyone’s opinion?
From a cost perspective, a closely related question is how many examples should be judged in the first place. A smaller test set not only lowers costs but also simplifies the task for any human reviewers — reading and evaluating 10 examples is far more manageable than 10,000.
If we assume that the main goal is to establish good relative rankings, we narrow the focus of quantifying the success of a Council by measuring the significance and stability of the final ranking. For significance, we look at separability (as in the main experiment), which measures how well models can be distinguished based on non-overlapping confidence intervals. For stability, we create a new metric, Mean Expected Rank Variance (MERV), defined as the expected ordinal swing of the average respondent’s rank (Appendix G). A MERV of 3 means an average respondent’s rank is expected to change up to 3 positions in a new trial. Lower MERV indicates that the benchmark has more stable rankings, with MERV of 0 signifying perfect deterministic-like stability.
@@ 5.1 Monte Carlo simulations
To study the dynamics of separability and stability with differently-sized councils and test sets irrespective of the inclusion of any specific LLM, we use a Monte Carlo procedure to simulate many random hypothetical councils and test sets. Our Monte Carlo simulation procedure is as follows:
  • (1)
    For a given council size $c$ and test set size $t$, randomly sample $c$ LLMs to form a council $C$ and $t$ examples to form a test set $T$. Sampling is performed with replacement.
  • (2)
    Find the associated judgments from the main experiment for the specific ($C$, $T$) configuration to determine the scores and relative rankings for all LLMs.
  • (3)
    Repeat for 100 trials. [^10] We use 100 to be consistent with the number of rounds of bootstrapping in the main experiment. For calculating separability in Monte Carlo simulations, the trials themselves are used to bootstrap confidence intervals directly.
  • (4)
    After all 100 trials are complete, tally the results: for stability, observe fluctuations in rankings for each LLM to compute MERV, and for significance, report the mean separability.
Figure 5 shows results for a sweep $c\in\{1,3,5,\dots,19\},\quad t\in\{10,20,30,\dots,100\}$.
@@
Both separability and stability improve as the number of test examples and the number of LLM judges increase, with the best scores achieved when both are maximized. However, based on the gradient maps for MERV (Figure 5(b)) and separability (Figure 5(d)), the added benefit of including either an additional judge or more test examples diminishes significantly in a concentric shape that starts ~50 examples and ~9 judges. The gradients in this darker zone indicate where the utility of any additional opinion—whether in the form of new test data or new LLM judges—becomes marginal.
@@
As the size of the LMC grows to include many different LLMs, it may become difficult to verify the quality of every member, particularly on subjective tasks. With the same Monte Carlo procedure, we experiment with a simulation setting with adversarial judges. An adversarial judge is a fake LLM council member that returns ratings at random. [^11] Other adversarial algorithms such as always voting for the first position were not explored. In Figure 6, we find that on both separability and stability, larger councils reduce the negative impact of adversarial LLM judges. This robustness continues to strengthen as the size of the council grows, even when maintaining the same ratio of adversarial judges to real judges, albeit with diminishing marginal returns.
@@
With respect to stability and separability, it depends. When test data is scarce, adding more test examples yields greater benefits than increasing the number of judges. However, once the test set exceeds 20-30 examples, introducing an additional judge becomes more valuable.
Larger councils also demonstrate greater resilience to adversarial judges. As the council size increases, the influence of any single unreliable judge diminishes, reducing the risk of significant disruption to the study. Consequently, strict selection criteria for council members become less critical in larger configurations.
@@ 5.2 Oligarchical councils
If a task does benefit from multiple perspectives and the evaluation budget allows for a limited number of opinions, whose opinions should be included? Can a subset of judges (or a single judge) effectively represent the fully democratic LMC?
We compare the rankings of three hand-curated sub-councils: flagships, smalls, and top-4.121212We refer to the sub-council as an “oligarchical” council because they rely on a small subset of LLM judges to determine rankings, akin to oligarchies in human society. See Table 9 for detailed sub-council membership.
While full council participation yields the highest scores for human agreement and statistical significance, our analysis finds that smaller sub-councils can still produce rankings aligned with human judgments while maintaining strong separability. Notably, smalls, a council composed of the smallest LLMs, achieves a separability of 71%—exceeding the average judge’s 53.3%—and a Spearman correlation of 0.88 with human rankings, only slightly below the full council’s 0.92 (Figure 23).
However, council composition remains a crucial factor. For instance, top-4 achieves the same correlation with human rankings as smalls (0.88) but with significantly higher separability (79%). While smaller councils can perform well, considerations such as separability, robustness, human alignment, and bias mitigation should be carefully weighed in council design.
@@ 6 Conclusion
In this paper, we introduce the Language Model Council (LMC), a flexible, decentralized evaluation framework for ranking LLM agents through democratic participation. Applying the LMC to an emotional intelligence task with 20 LLMs, we demonstrate that the LMC can produce highly separable rankings that align more closely with human judgments than other benchmarks or individual judges. Through both Monte Carlo simulations and hand-curated sub-councils, we find that while larger councils provide benefits, they are incrementally diminishing, and the majority of key qualities—such as ranking significance, stability, and alignment with human evaluations—can be achieved with a smaller, well-chosen ensemble of judges. As humans increasingly rely on LLMs to evaluate other LLMs, we hope the LMC framework, along with insights from our case study, offers a valuable foundation for developing reliable yet inclusive LLM evaluations, even for highly subjective tasks.
@@
LLMs are inherently stochastic, meaning the same model can produce different ratings even with temperature set to 0 Chann (2023). Reproducibility is further complicated by closed-weight models like GPT-4, which may receive undisclosed updates. All responses in our study were collected in May 2024, but serverless providers, like Together [^13] https://www.together.ai/, may update their APIs or discontinue support for certain models, as happened with Qwen-1.5-32B (replaced by Qwen-2.5). For open source models, changes in deployment hardware or GPU configuration can introduce additional variability, making exact replication of results difficult.
@@
While the LMC effectively neutralizes biases individual LLMs, it is still not immune to systemic biases within the evaluation framework itself. Our case study suggests that using a single reference model may have inadvertently favored its successors within the arena. To mitigate such biases, we recommend conducting dry runs with individual LLM judges to detect and correct mechanical biases in evaluation design before scaling to a full council.
@@ Acknowledgements
We thank Sahand Sabour for creating EmoBench and insightful discussions about emotionally rich synthetic data. We thank Alex Tamkin for proposing the idea to measure the value of the incremental judge. We thank Mitchell Gordon for his suggestion to qualitatively analyze why certain responses are preferred. We thank David So for his ideas on calibration, repeatability, and model affinity. We thank Federico Bianchi and Sam Paech for their idea of measuring oligarchical councils and for their feedback on prompt design. Sam Paech provided insightful discussions on separability, voting aggregation, length bias, and the relationship to other leaderboards, and he advised us throughout the project. Finally, we thank Predibase for their support and funding this research.
Flor Miriam Plaza-del-Arco is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 949944, INTEGRATOR). She is a member of the MilaNLP group and the Data and Marketing Insights Unit of the Bocconi Institute for Data Science and Analysis (BIDSA). Amanda Cercas Curry is a former member of MilaNLP and was supported by the same grant while working on this study.
Appendix A Additional Findings
@@
A note on proactively remove inconsistent LLM judge ratings
In the two-game setup during pairwise comparisons, we gather ratings for models in both positions, allowing us to identify and potentially remove inconsistent ratings before calculating ELO scores.
Automatically removing inconsistent votes affects the weighting of LLM judges, as those with more inconsistent ratings will have fewer votes counted and thus less influence overall. However, this could also be argued as a positive outcome, as positional consistency is often a sign that the judgment was noisy or arbitrary, especially when tie ratings are not permitted. By comparison, some arena-based evaluation systems like Zheng et al. (2023) allow judges to declare a tie between responses, which are subsequently excluded from ELO scoring.
In our EI case study, we chose to retain all inconsistent votes, allowing downstream processes like aggregation, Bradley-Terry scoring, and bootstrapping to promulgate any diminished influence caused by inconsistent voting.
Since inconsistent votes are a potential source of noise, however, we include metrics when only consistent votes are considered. For the hand-selected sub-councils analyzed in Section 5.2, Table 3 shows the changes in key judging qualities when considering only consistent votes. Table 4 shows the impact of inconsistent vote pre-filtering on length bias.
@@
The 20x20 LLM interactions generate a wealth of data that spans multiple pages. For ease of reference, all large tables and figures are compiled here:
@@
Figure 10: Heatmap of the affinities of LLM judges to LLM respondents. The relatively consistent horizontal bands in the heatmap suggest a clear consensus on the preferred LLM participants. Some judges, like mistral-large, exhibit high polarization, with a significant difference between their highest and lowest-rated LLMs. In contrast, judges like llama-3-8b display a narrower range of affinity. The highest affinity expressed by any LLM comes from mistral-large for Qwen-1.5-110B, while the strong horizontal blue band for gemini-1.5-pro indicates it was a consensus low performer.
@@
Figure 12: Graph of top 5 affinities. An edge exists from LLM $a$ to LLM $b$ if $\text{affinity}(a,b)$ is in the top 5 affinities for LLM $a$. This view allows us to identify "popular" LLMs, "hipster" LLMs, and "LLM friendships" (where two LLMs have strong mutual affinity for each other). Popular LLMs, such as GPT-4o, have many arrows pointing toward them, while unpopular LLMs like dbrx-instruct have none. Qwen-1.5-32B and command-r form a "friendship" with mutual strong affinities, though Qwen-1.5-32B also receives incoming edges from three other LLMs, making it more widely liked. Some LLMs have only one "fan," others have several, and some have none. Of course, this analysis is somewhat contrived, as affinities are continuous values and using the top 5 as a cutoff is arbitrary. Nevertheless, it offers an interesting starting point for studying the dynamics and patterns that emerge when an arbitrary affinity threshold is established.
@@
United States Open AI gpt-4-0613 OpenAI (2023) 06/23 1246 86.4 [Uncaptioned image] Proprietary
United States Meta llama-3-70b-chat-hf Platforms (2024) 04/24 1208 82.0 70 B Llama 3 Community
United States Meta llama-3-8b-chat-hf Platforms (2024) 04/24 1153 68.4 8 B Llama 3 Community
Canada Cohere command-r Cohere (2024a) 04/24 1147 68.2 35 B CC-BY-NC-4.0
China Alibaba qwen1.5-110B-chat Team (2023) 02/24 1164 80.2 100 B Qianwen LICENSE
China Alibaba qwen1.5-72B-chat Team (2023) 02/24 1152 77.4 72 B Qianwen LICENSE
China Alibaba qwen1.5-32B-chat Team (2023) 02/24 1126 74.3 32 B Qianwen LICENSE
LLM All Flagships Smalls Top-4
gpt-4-turbo-04-09 [Uncaptioned image] [Uncaptioned image]
gpt-4-0613 [Uncaptioned image]
gpt-3.5-turbo-0125 [Uncaptioned image]
mistral-large-latest [Uncaptioned image] [Uncaptioned image]
open-mixtral-8x22b [Uncaptioned image]
open-mixtral-8x7b [Uncaptioned image] [Uncaptioned image]
llama-3-70b-chat-hf [Uncaptioned image] [Uncaptioned image]
llama-3-8b-chat-hf [Uncaptioned image] [Uncaptioned image]
gemini-1.0-pro [Uncaptioned image] [Uncaptioned image]
dbrx [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
command-r-plus [Uncaptioned image] [Uncaptioned image]
command-r [Uncaptioned image] [Uncaptioned image]
claude-3-sonnet-20240229 [Uncaptioned image]
claude-3-haiku-20240307 [Uncaptioned image] [Uncaptioned image]
qwen1.5-110B-chat [Uncaptioned image] [Uncaptioned image]
qwen1.5-72B-chat [Uncaptioned image]
qwen1.5-32B-chat [Uncaptioned image] [Uncaptioned image]
Appendix B LLM Judge Calibration
To understand the reliability and natural variability of LLM model judges and to help us decide evaluation settings, we run a calibration exercise prior to the main experiment.
We collect pairwise preference ratings on three responses to the same interpersonal conflict using different temperatures and pairwise comparison options. Two responses are competitive, and one is intentionally generic to serve as a ranking baseline (Figure 17). We measure Invariability and Pairwise Positional Consistency (PPC), defined below.
B.1 Invariability